40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6896     Entry time: Fri Jun 29 16:41:41 2012
Author: yuta 
Type: Update 
Category: General 
Subject: PRM was NOT installed backwards 

[Koji, Steve, Jamie, Yuta]

So, PRM was NOT flipped......

We opened the BS chamber and quickly checked the arrow on the PRM pointing HR. It turned out to be correct, the arrow was pointing towards the arm cavity. We opened the ITMX chamber, too, to check PR2 later.
BS chamber and ITMX chamber is now closed with the light door.

But it was a one step forward anyway, because we could prove PRM was innocent.

What to do next:
  We know that the mode-matching of the incident beam and both arms are pretty good. So, dirty modes come from PRC.
  We will check beam clipping, mirrors, suspensions in PRC.
  I expect the chambers to be closed on Monday(July 2) afternoon and start pumping on Tuesday(July 3) morning.

ELOG V3.1.3-